A compound is composed of 23.0 g sodium and 35.0 g chlorine is available for reaction, what mass of sodium is needed to complete the salt reaction?
1
Expert's answer
2020-10-20T08:34:17-0400

Explanation:

The reaction that takes place is:

Na + Cl -> NaCl

the ratio of this reaction is 1 : 1 : 1

Na =22.9898 g/mole g/mole

Cl = 35.453 g/mole

NaCl = 58.44 g/mole

First you calculate the mole using the known data:

mole = mass / molar mass

mole sodium = 23.0 / 22.9898 = 1.00

since the ratio is 1 : 1 : 1, all you have to do is:

mass = mole * molar mass

mass NaCl = 1.00 * 58.44 = 58.44 g

you can check if all the mass of Cl is used by calculating the mass using the mole and the molar mass:

Mass Cl = 1.00 * 35.0 = 35.0 g

Which was the amount of mass used in the reaction so everything correct, meaning the amount of mass formed in the reaction is 58.44 grams.